OpenStudy (anonymous):

A moving company charges a base rate of $55.00 plus $20.00 per hour. If the moving company charged Mr. Sanchez $110.00, which equation could be used to determine the number of hours it took the moving company? (h = hour)

OpenStudy (wach):

Do you have any ideas, knowing that the move costs at least 55$ plus 20$ dependent on the variable of h, hours?

OpenStudy (anonymous):

not really good at math.!!!

OpenStudy (wach):

I'm trying to help you work though it. So let's look at what we know: The base rate is $55, so that means that 55 of the 110 dollars Mr Sanchez spent was not affected by time. This means that the other portion of money he spent (110-55=55), 55 dollars was dependent on time.

OpenStudy (wach):

If he spend $55 dollars on x amt of hours, each costing $20, how much did he spend?

OpenStudy (wach):

Try dividing 55$ by the cost per hour, 20$, to get the total amount of hours.

OpenStudy (wach):

Did I lose you?

OpenStudy (anonymous):

charges a base rate of $55.00 plus $20.00 per hour $110 = $55 + $20 h => h = (110 - 55 ) / 20 = ...